OpenStudy (anonymous):

Given the function f(x) = 6(x+2) − 3, solve for the inverse function when x = 21.

OpenStudy (zpupster):

lets review the steps of finding an inverse function: Step 1: Replace f(x) with y. Step 2: Interchange x with y. Step 3: Solve for y. If this y is a function, it is the inverse of the original function. If this y is NOT a function, there is no inverse. Step 4: If f has an inverse function, then replace y with f inverse. ok lets walk thru this problem first lets rearrange the problem. Given the function f(x) = 6(x+2) − 3 f(x)= 6x+12 -3 = 6x+9 Step 1 y=6x+9 step 2 x=6y+9 step3 can you solve for y?

OpenStudy (anonymous):

thank you